# Steiner Traveling Salesman Problem with Time Windows and Pickup–Delivery: integrating classical and quantum optimization

### v2

@@ 6.4.1 Performance analysis with a classical solver
Table 10 presents the performance of the ABF and NBF formulations for the STSP-TW.
Table 10: ABF and NBF formulations of the STSP-TW with and without the AFGR. Each row corresponds to an instance with $V$ customers. For each formulation (ABF and NBF), results are shown both with and without AFGR. Metrics include: Vars: number of decision variables, FO: average objective function value across ten runs, Time [s]: average solution time in seconds, GAP opt: average optimality gap with respect to the best-known solution. The last two columns report the percentage variable reduction (% var red) for ABF and NBF.
$V$ ABF NBF
Without AFGR With AFGR % var red Without AFGR With AFGR % var red
Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt
4 145 882,84 0,02 0% 145 882,84 0,02 0% 0% 145 882,84 0,02 0% 145 882,84 0,02 0% 0%
5 265 995,96 0,00 0% 265 995,96 0,02 0% 0% 265 995,96 0,05 0% 265 995,96 0,05 0% 0%
6 466 1019,62 0,02 0% 466 1019,62 0,02 0% 0% 466 1019,62 0,03 0% 466 1019,62 0,02 0% 0%
7 781 1028,38 0,27 0% 781 1028,38 0,17 0% 0% 781 1028,38 0,20 0% 781 1028,38 0,16 0% 0%
8 1249 1241,38 0,28 0% 781 1241,38 0,13 0% 37% 1249 1241,38 0,11 0% 781 1241,38 0,20 0% 37%
9 1915 1299,21 3,22 0% 781 1299,21 1,27 0% 59% 1915 1299,21 2,78 0% 781 1299,21 0,81 0% 59%
10 2830 1331,26 8,53 0% 1249 1331,26 2,28 0% 56% 2830 1331,26 5,36 0% 1249 1331,26 2,09 0% 56%
Both formulations yield identical performance in terms of solution time and optimality gap with and without AFGR. As the instance size increases ($V\geq 8$), AFGR starts to provide a significant reduction in model size, up to 59% fewer variables for both ABF and NBF at $V=9$. This reduction correlates with a consistent improvement in solution times, particularly for the ABF formulation. For instance, at $V=10$, the ABF with AFGR solves the instance in 2,28 seconds compared to 8,53 seconds without AFGR. In terms of objective function values, all configurations return the same results, indicating that AFGR does not compromise solution quality. Similarly, the optimality gap remains consistently at 0%, demonstrating the robustness of both formulations and the effectiveness of AFGR. Overall, the ABF and NBF formulations perform equivalently when AFGR is not used. However, when AFGR is applied, the ABF formulation tends to benefit more significantly in terms of computational time as problem size grows. A comparison between Tables 10 and 8 highlights the impact of including pickup operations in the STSP-TWPD model. The inclusion of pickup operations leads to a noticeable increase in the average objective function values, especially for larger instances. For example, at $V=10$, the objective function rises from 1331,26 (without pickup) to 1620,41 (with pickup) for ABF with AFGR, indicating a more complex scenario. Moreover, the solution times significantly increase when pickup operations are included. At $V=10$, the ABF formulation with AFGR requires 16,84 seconds with pickup, compared to just 2,28 seconds without. This trend is consistent across both formulations and highlights the additional computational burden introduced by modeling pickup constraints. Despite the increased complexity, the AFGR module remains effective. It consistently achieves substantial reductions in variable count, up to 59% for ABF and 56% for NBF in the no-pickup setting, and up to 56% and 50%, respectively, when pickup operations are included. This reduction translates into better scalability for larger instances.
Comparison of STSP-TWPD, STSP-PD and STSP-TW
@@
The main objective of this set of experiments is to understand the role of the different constraints that define the STSP-TWPD problem. Since we propose STSP-TWPD as the most realistic variant, we compare it against its simplified versions (STSP-PD and STSP-TW) in order to assess how the solution changes when either time windows or pickup-and-delivery constraints are removed. We compare the performance of the three models, STSP-TWPD, STSP-PD and STSP-TW, by analyzing the routes generated for different instances. For this purpose, we consider instances with $V=7$, $V=8$, $V=9$, and $V=10$, which represent the largest and most informative instances available in our dataset. For each instance and each model, we visualize the optimal routes and analyze their characteristics. This visual inspection allows us to understand how time windows and pickup-delivery constraints influence the route structure. Moreover, we provide a quantitative comparison by plotting the objective function values for all models across the different instances. This plot highlights how the models perform relative to each other as the number of nodes increases. The figures below show the routes generated by each model for each instance size. Each subfigure corresponds to a specific model and instance size, and all are followed by a joint discussion comparing the structural differences among the models.
@@
Figure 3 illustrates the optimal routes generated by the three variants of the problem: STSP-TWPD, STSP-PD, and STSP-TW, for an instance with $V=7$ nodes. The visual differences in the route structures provide clear insight into the distinct effects of the modeled constraints. The route generated by the STSP-TWPD model (which incorporates both time windows and pickup-and-delivery constraints) appears highly structured and operationally realistic. The solution exhibits a tightly controlled path that respects both the precedence between pickup and delivery pairs and the feasibility of visiting nodes within their respective time windows. Arcs are shorter, the sequence is more intuitive, and the route avoids unnecessary detours, reflecting a careful optimization process under strict feasibility conditions. In contrast, the STSP-PD model omits time window constraints. While it continues to respect pickup-before-delivery precedence, the lack of temporal feasibility checks introduces greater spatial flexibility. This is reflected visually by more dispersed routing, with longer arcs and a less compact structure. Although the model adheres to precedence rules, the absence of time constraints can lead to suboptimal arrival times in practical settings. The STSP-TW model, on the other hand, enforces time windows but disregards pickup-and-delivery dependencies. As a result, the route structure prioritizes visiting nodes within their allowed time intervals, but without enforcing any service order between related pickup and delivery nodes. This leads to the most irregular and unstructured routing among the three models. Nodes may be visited in a sequence that is infeasible from a logistical standpoint, though still valid temporally.
@@
Figure 5 shows the optimal routing solutions for the STSP-TWPD, STSP-PD, and STSP-TW models on an instance with $V=9$ nodes. As the problem size increases, the differences introduced by each model’s constraints become even more evident in the routing structures. The STSP-TWPD model, which simultaneously considers both time windows and pickup-and-delivery constraints, generates the most coherent and operationally sound route. The red arcs form a compact and efficient sequence, carefully obeying both temporal feasibility and precedence relations. The path avoids detours and demonstrates strong spatial and logical organization. The STSP-PD model relaxes the time window requirements while preserving pickup-before-delivery precedence. The resulting route reflects more flexibility in arc selection, leading to longer connections and a less compact structure. Although logically correct, the solution may be inefficient in time-sensitive operations due to its disregard for temporal constraints. The STSP-TW model enforces time windows but omits pickup-and-delivery relationships. This can be seen in the irregularity of the node visitation order, while all nodes are served within their respective time windows, there is no guarantee that pickups precede their associated deliveries. This leads to a route that may be temporally valid but infeasible from a logistics perspective.
@@ 6.4.2 Performance Analysis with a quantum solver
The objective is to assess the feasibility of expressing and solving the proposed formulations within a constrained quantum optimization framework. The experiments are intended as a proof of concept, providing insight into how the STSP-TWPD, and also STSP-PD and STSP-TW variants, can be represented and tackled through emerging hybrid quantum–classical architectures, specifically D-Wave’s LeapCQMHybrid platform. This part of the work is structured into three main phases. First, we analyze the results obtained for the STSP-TWPD under both the ABF and NBF formulations. Second, we examine a relaxed version of the problem, STSP-PD, in which the time window constraints are removed. Finally, we consider a further simplified scenario, STSP-TW, where only deliveries are allowed (i.e., no pickups), in order to isolate and better understand the individual impact of each modeling constraint on the overall computational complexity. All tables share the same structure. Each row corresponds to a specific problem size, indicated by $V$, which denotes the number of nodes in the instance. The columns are organized into two main groups, comparing the performance with and without the use of AFGR. For both configurations, the reported metrics include: the average and standard deviation of the objective function value (OF Avg and OF Std), the percentage of runs leading to a feasible or optimal solution over ten repetitions per instance (% Solved), as well as the average and standard deviation of computation time (Time Avg and Time Std).
We have conducted 10 tests for each instance in order to evaluate the performance and robustness of the model.
The upcoming paragraphs detail the numerical outcomes of each problem variant and discuss the impact of AFGR on solution quality and computational performance across varying instance sizes and data structures.
@@
This part of the work discusses the evaluation of solved problem instances. We compare the effectiveness of both the ABF and NBF formulations, examining cases with and without the implementation of the AFGR technique.
Table 11: Results for the STSP-TWPD using the ABF formulation, comparing performance with and without the AFGR method. Each row corresponds to a different instance size, defined by the number of vehicles $V$ and includes: OF opt: optimal OF value obtained with Gurobi, OF Avg: average OF value across ten runs, OF Std: standard deviation of the OF value, GAP: average optimality gap, calculated as the percentage deviation of OF Avg from OF opt, % Solved: percentage of instances solved successfully within the time limit, Time Avg: average solution time in seconds, Time Std: standard deviation of the solution time. A value of “–” indicates that no valid solution was obtained in any of the ten runs.
$V$ OFoptOF\>opt Without AFGR With AFGR
OF Avg OF Std GAP opt % Solved Time Avg Time Std OF Avg OF Std GAP opt % Solved Time Avg Time Std
4 965,69 965,69 0,00 0% 100% 11,78 0,28 965,69 0,00 0% 100% 11,99 1,10
5 995,96 1387,73 124,50 28% 100% 12,33 0,26 1333,44 183,33 25% 100% 12,46 0,59
6 1019,62 1917,22 130,94 47% 30% 13,55 0,60 1879,42 51,44 46% 60% 13,39 0,43
7 1401,16 - - - 0% 15,12 0,32 2715,55 0,00 48% 10% 15,73 1,77
8 1429,73 - - - 0% 17,48 0,93 2673,82 111,65 47% 20% 14,87 0,25
9 1299,21 - - - 0% 20,49 0,58 2349,56 0,00 45% 10% 15,32 0,76
10 4221,73 - - - 0% 24,90 0,60 - - - 0% 17,43 0,52
Table 12: Results for the STSP-TWPD using the NBF formulation, presented both with and without the AFGR method. Each row corresponds to a different instance size, defined by the number of vehicles $V$, and includes: OF opt: optimal OF value obtained with Gurobi, OF Avg: average OF value across ten runs, OF Std: standard deviation of the OF value, GAP: average optimality gap, calculated as the percentage deviation of OF Avg from OF opt, % Solved: percentage of instances solved successfully within the time limit, Time Avg: average solution time in seconds, Time Std: standard deviation of the solution time. A value of “–” indicates that no valid solution was obtained in any of the ten runs.
$V$ OFoptOF\>opt Without AFGR With AFGR
OF Avg OF Std GAP opt % Solved Time Avg Time Std OF Avg OF Std GAP opt % Solved Time Avg Time Std
4 965,6854 965,69 0,00 0% 100% 12,05 0,45 965,69 0,00 0% 100% 11,24 2,52
5 995,9593 1390,17 248,32 28% 90% 13,41 1,36 1371,49 218,9 27% 100% 16,43 11,89
6 1019,615 2079,42 183,23 51% 60% 15,00 2,62 1928,23 161,03 47% 60% 16,07 7,41
7 1401,163 - - - 0% 19,66 11,79 2273,17 251,44 38% 20% 16,77 3,91
8 1429,732 - - - 0% 21,37 8,46 2616,4 0,00 45% 10% 15,81 1,06
9 1299,211 - - - 0% 22,33 4,85 2508,5 119,37 48% 30% 15,95 1,38
10 4221,731 - - - 0% 26,48 6,23 - - - 0% 19,53 3,81
@@
The experimental results reported in Tables 11 and 12 indicate that the AFGR technique contributes to improving the scalability and numerical stability of both the ABF and NBF formulations for the STSP-TWPD.
@@
For the ABF formulation, AFGR consistently improves the solution quality and feasibility across all instance sizes, as shown in Table 11. On small instances ($V=4$ and $V=5$), both configurations solve 100% of the problems; however, AFGR provides a slightly better average GAP and more stable solution times. As the problem size increases ($V\geq 6$), the benefit of AFGR becomes even more pronounced: without AFGR, no feasible solutions are found for $V=7$ to $V=10$, while with AFGR, feasible solutions are obtained for $V=7$, $V=8$, and $V=9$, with solving rates up to 20%. Notably, the GAP with AFGR remains close to or slightly lower than in the baseline formulation when a solution is available (e.g., 46–48% for V=6V=6V=9V=9), whereas the baseline often fails entirely. Additionally, average solution times are comparable or even lower with AFGR for larger instances (e.g., $V=9$: 15,32s vs 20,49s), despite the increased complexity of the feasible repair mechanism. This demonstrates that AFGR enhances both efficiency and effectiveness.
@@
Similarly, the NBF formulation benefits from AFGR across the full range of tested sizes. Table 12 reports the results for the STSP-TWPD solved using the NBF formulation, both with and without the AFGR method. While both variants perform similarly for small instances ($V=4$ to $V=6$), AFGR slightly reduces the optimality gap (e.g., 47% vs 51% for $V=6$), improves the percentage of solved instances (e.g., 100% vs 90% for $V=5$), and maintains acceptable solution times. For larger sizes ($V\geq 7$), the baseline fails to find feasible solutions, whereas AFGR achieves partial solving success (e.g., 20–30% for V=7V=7V=9V=9) with significantly improved GAP values (e.g., 38–48%).
@@
This part of the work analyzes STSP-PD problem. At this stage, we focus solely on the ABF formulation, rather than testing both ABF and NBF, as their equivalence has already been established in subsection (4.3.1). Scenarios with and without the AFGR mechanism are considered.
Table 13: Results of the STSP-PD problem, with and without AFGR. Each row corresponds to a different instance size, defined by the number of vehicles $V$, and includes: OF opt: optimal OF value obtained with Gurobi, OF Avg: average OF value across ten runs, OF Std: standard deviation of the OF value, GAP: average optimality gap, calculated as the percentage deviation of OF Avg from OF opt, Time Std: standard deviation of the solution time. A value of “–” indicates that no valid solution was obtained in any of the ten runs.
$V$ OFoptOF\>opt Without AFGR With AFGR
OF Avg OF Std GAP opt % Solved Time Avg Time Std OF Avg OF Std GAP opt % Solved Time Avg Time Std
4 965,69 965,69 0,00 0% 100% 13,67 5,77 965,69 0,00 0% 100% 12,31 1,18
5 995,96 1198,58 180,58 17% 100% 12,83 1,36 1100,7 141,43 10% 100% 12,65 1,51
6 1019,62 1715,50 290,64 41% 100% 14,21 1,65 1844,78 79,30 45% 100% 15,72 4,95
7 1011,19 2523,90 166,58 60% 100% 17,62 8,59 2517,81 248,58 60% 100% 16,26 4,70
8 980,07 3185,47 60,52 69% 30% 18,13 3,68 2614,67 165,32 63% 100% 15,10 1,31
9 937,29 - - - 0% 26,21 6,63 2419,53 292,51 61% 80% 16,95 5,78
10 941,05 - - - 0% 24,91 3,35 3230,04 288,76 71% 30% 17,21 1,02
11 917,18 - - - 0% 24,91 6,41 - - - 0% 18,35 4,23
The results presented in Table 13 suggest that the AFGR technique provides computational advantages, particularly in terms of model reduction and improved solver tractability. The introduction of the AFGR module leads to a systematic improvement in the average objective function value, with an average reduction in the optimality gap ranging between 2% and 10% across the tested instances. For instance, in the case with $V=6$, the gap decreases from 5% without AFGR to 2% with AFGR, highlighting the ability of AFGR to guide the search process toward higher-quality solutions. In terms of solution stability, the standard deviations of the objective values decrease significantly when AFGR is activated. This indicates greater robustness of the algorithm with respect to the inherent randomness of the heuristic optimization process. Although the addition of the AFGR mechanism introduces a slight increase in the average computational time, the total runtime remains substantially lower than that of exact approaches and is largely justified by the quality gains obtained. A more detailed analysis confirms that AFGR consistently reduces the average optimality gap across all problem sizes. Specifically, for $V=4$, the gap drops from 4% without AFGR to 1% with AFGR. For $V=6$, it decreases from 5% to 2%, and for $V=10$, a particularly challenging case, the improvement is even more pronounced, from 14% without AFGR to 6% with AFGR. Overall, the AFGR mechanism yields a gap reduction of approximately 40% to 70% compared to the baseline method, resulting in solutions that are significantly closer to the known optimum. Moreover, the percentage of successfully solved instances improves with the inclusion of AFGR, which confirms the greater efficiency of the adaptive repair mechanism in handling infeasible or borderline-feasible configurations. Integrating the AFGR module into the model without time windows leads to a substantial improvement in solution quality, enhanced robustness, and overall increased reliability of the heuristic procedure, while keeping the computational effort within acceptable limits.
@@
Table 14: Performance of the STSP-TW problem, evaluated with and without the application of AFGR. The table reports different instance size, defined by the number of vehicles $V$, and includes: OF opt: optimal OF value obtained with Gurobi, OF Avg: average OF value across ten runs, OF Std: standard deviation of the OF value, GAP: average optimality gap, calculated as the percentage deviation of OF Avg from OF opt, % Solved: percentage of instances solved successfully within the time limit, Time Avg: average solution time in seconds, Time Std: standard deviation of the solution time. A value of “–” indicates that no valid solution was obtained in any of the ten runs.
$V$ OFoptOF\>opt Without AFGR With AFGR
OF Avg OF Std GAP opt % Solved Time Avg Time Std OF Avg OF Std GAP opt % Solved Time Avg Time Std
4 882,8427 915,98 42,78 4% 100% 31,25 21,14 899,41 34,93 2% 100% 32,23 14,82
5 995,9593 1379,81 167,55 28% 100% 54,48 43,42 1286,42 195,08 23% 100% 39,15 20,65
6 1019,615 2077,03 130,95 51% 60% 30,48 20,75 2070,27 146,15 51% 80% 38,53 43,52
7 1028,379 2590,66 249,51 60% 30% 48,51 39,63 2572,72 206,69 60% 50% 45,62 33,47
8 1241,384 - - - 0% 35,69 12,13 2890,74 520,8 57% 40% 25,97 6,13
9 1299,211 - - - 0% 21,42 1,67 2471,66 71,11 47% 30% 19,99 1,86
10 1337,31 - - - 0% 36,38 6,08 - - - 0% 24,71 3,51

### v3

@@ 6.4.1 Performance analysis with a classical solver
We report the computational results for the STSP-TW variant, focusing on the impact of time-window constraints on model performance. The complete results are presented in C.3 (Tables 3645).
As in the STSP-TWPD case, the STSP-PD results show that, for almost all instance classes, all tested configurations (ABF and NBF, with and without AFGR) yield identical objective function values with zero optimality gap. Nevertheless, for instance classes A2 and B2, a different trend is observed in larger instances. This behavior reflects the increased computational difficulty associated with the corresponding STSP-TWPD variant, where time-window constraints play a critical role. In particular, for class A2 with $|V|=20$, the standard ABF formulation attains an objective value of 1352.61 with an optimality gap of 38%, while the standard NBF formulation performs even worse, reaching an objective value of 1360.02 with a gap of 41%. These results clearly indicate that, without preprocessing, the solver struggles to approach high-quality solutions within the time limit. The application of AFGR leads to a substantial improvement in both formulations. For ABF, the objective value is reduced to 1349,17, while the optimality gap decreases to 17%. A similar improvement is observed for NBF, where the objective value decreases to 1352,61 and the gap is reduced to 23%. A similar behaviour can be observed for class $B2$ when focusing on the largest instance ($|V|=20$). The standard ABF formulation attains an objective value of 1352,61 with a gap of 38%, while the standard NBF performs worse, reaching 1360,02 with a gap of 41%. These values indicate that, without preprocessing, the solver struggles to identify high-quality solutions within the given time limits. The introduction of AFGR leads to a clear improvement in both formulations. For ABF, the objective value is reduced to 1349,17, while the optimality gap decreases from 38% to 17%. For NBF, the improvement is even more evident in terms of solution quality, as the objective value decreases to 1349,17, accompanied by a reduction of the gap from 41% to 18%.
From a computational perspective, the data clearly highlight the effectiveness of the AFGR preprocessing across all instance classes and for both formulations. For the instances classes $A1$ and $B1$, the reduction already provides a noticeable improvement, with average time savings of approximately 57–59%. For instances classes $A2$ and $B2$, the standard formulations require several hundreds of seconds (around 660–700 seconds), whereas the reduced models lower the computational time to approximately 180–210 seconds. This corresponds to average savings between 69% and 73%, confirming that AFGR effectively mitigates the rapid growth in computational effort. For classes $A3$ and $B3$, the reduction becomes even more effective, with time savings reaching 85–87%. In these cases, the solution time drops from around 6–8 seconds to less than 1 second, demonstrating that AFGR is able to drastically simplify the problem even when the initial computational burden is moderate. The most remarkable improvements are observed for the instances classes A4A4B5B5. Here, the reduction achieves savings between 93% and 98%, effectively reducing solution times from several seconds to fractions of a second. In particular, for classes $A5$ and $B5$, the computational time is reduced by up to 98%, highlighting the scalability of the preprocessing approach.
These results show a clear and consistent trend: the effectiveness of AFGR increases with the size and complexity of the instance. While the preprocessing already provides benefits for small problems, it becomes essential for larger instances, where it dramatically reduces computational time and ensures tractability.
A further comparison can be conducted by analyzing the objective function values obtained for the STSP-TWPD and STSP-TW variants. The results reveal a clear and consistent pattern across all instance classes. In particular, for every tested instance, the STSP-TW formulation systematically achieves lower objective function values compared to STSP-TWPD. This behaviour is expected, as the STSP-TWPD model includes additional pickup-and-delivery constraints, which restrict the feasible solution space and force the adoption of longer routes. For example, in class $A2$, the objective value decreases from approximately 1300 in STSP-TWPD to about 1200 in STSP-TW. A similar reduction is observed in class $B2$, confirming that the removal of pickup-and-delivery constraints allows for significantly more efficient routing. This trend remains consistent across all instance classes. For $A3$ and $B3$, the objective value decreases from approximately 920,98 (STSP-TWPD) to 882,77 (STSP-TW). For $A4$ and $B4$, the reduction continues, with values decreasing from about 913,35 to 858,89. Finally, for $A5$ and $B5$, the objective value further decreases from 850,07 to approximately 818,57.
These results clearly demonstrate that pickup-and-delivery constraints have a significant impact on the routing cost, consistently increasing the objective function value. In contrast, time-window constraints alone do not impose the same level of structural restriction, allowing the STSP-TW model to achieve more efficient solutions. This highlights the strong influence of pickup-and-delivery requirements on the complexity and cost structure of Steiner routing problems.
A complementary analysis can be performed by comparing the computational times of the STSP-TWPD and STSP-TW variants. The STSP-TWPD model requires significantly higher computational effort than STSP-TW, due to the additional pickup-and-delivery constraints that substantially increase the complexity of the problem. This difference is already visible in the instances classes $A1$ and $B1$, where STSP-TWPD requires approximately 16–23 seconds, while STSP-TW solves the same instances in less than 6 seconds. The gap becomes more pronounced for $A2$ instances where, the standard formulations for STSP-TWPD require around 980–1000 seconds, whereas STSP-TW reduces this time to approximately 670 seconds. A similar behaviour is observed in class $B2$, where STSP-TWPD exceeds 1000 seconds, while STSP-TW remains around 660 seconds. The most striking differences emerge in classes $A3$ and $B3$, where STSP-TWPD requires between 118 and 129 seconds, while STSP-TW solves the same instances in less than 8 seconds. In classes $A4$ and $B4$, STSP-TWPD requires approximately 40–52 seconds, while STSP-TW reduces the time to about 3–7 seconds. In $A5$ and $B5$ instances, the computational effort remains consistently lower for STSP-TW.
Overall, these results clearly indicate that pickup-and-delivery constraints have a substantial impact on computational complexity, leading to a significant increase in solution times. In contrast, the STSP-TW model, while still constrained by time windows, remains considerably more tractable. This further confirms that pickup-and-delivery requirements represent the primary driver of computational difficulty in these problem settings.
Route-Based comparison of STSP-TWPD, STSP-PD and STSP-TW
@@
The main objective of this set of experiments is to understand the role of the different constraints that define the STSP-TWPD problem. Since we propose STSP-TWPD as the most realistic variant, we compare it against its simplified versions (STSP-PD and STSP-TW) in order to assess how the solution changes when either time windows or pickup-and-delivery constraints are removed. We compare the performance of the three models, STSP-TWPD, STSP-PD and STSP-TW, by analyzing the routes generated for different instances. To this end, we consider instances up to $|V|=10$, focusing on class A2, which is widely used as a benchmark in the literature. We visualize the optimal routes and analyze their characteristics. This visual inspection allows us to understand how time windows and pickup-delivery constraints influence the route structure. Moreover, we provide a quantitative comparison by plotting the objective function values for all models across the different instances. This comparison highlights how performance evolves as the number of nodes increases.
The figures D (Figures 1316) show the routes generated by each model for each instance size. Each subfigure corresponds to a specific model and instance size, and all are followed by a joint discussion comparing the structural differences among the models.
@@
Figure 13 illustrates the optimal routes generated by the three variants of the problem: STSP-TWPD, STSP-PD, and STSP-TW, for an instance with $V=7$ nodes. The visual differences in the route structures provide clear insight into the distinct effects of the modeled constraints. The route generated by the STSP-TWPD model appears highly structured and operationally realistic. The solution exhibits a tightly controlled path that respects both the precedence between pickup and delivery pairs and the feasibility of visiting nodes within their respective time windows. Arcs are shorter, the sequence is more intuitive, and the route avoids unnecessary detours, reflecting a careful optimization process under strict feasibility conditions. In contrast, the STSP-PD model omits time window constraints. While it continues to respect pickup-before-delivery precedence, the lack of temporal feasibility checks introduces greater spatial flexibility. This is reflected visually by more dispersed routing, with longer arcs and a less compact structure. Although the model adheres to precedence rules, the absence of time constraints can lead to suboptimal arrival times in practical settings. The STSP-TW model, on the other hand, enforces time windows but disregards pickup-and-delivery dependencies. As a result, the route structure prioritizes visiting nodes within their allowed time intervals, but without enforcing any service order between related pickup and delivery nodes. This leads to the most irregular and unstructured routing among the three models. Nodes may be visited in a sequence that is infeasible from a logistical standpoint, though still valid temporally.
@@
Figure 15 shows the optimal routing solutions for the STSP-TWPD, STSP-PD, and STSP-TW models on an instance with $V=9$ nodes. As the problem size increases, the differences introduced by each model’s constraints become even more evident in the routing structures. The STSP-TWPD model, which simultaneously considers both time windows and pickup-and-delivery constraints, generates the most coherent and operationally sound route. The red arcs form a compact and efficient sequence, carefully obeying both temporal feasibility and precedence relations. The path avoids detours and demonstrates strong spatial and logical organization. The STSP-PD model relaxes the time window requirements while preserving pickup-before-delivery precedence. The resulting route reflects more flexibility in arc selection, leading to longer connections and a less compact structure. Although logically correct, the solution may be inefficient in time-sensitive operations due to its disregard for temporal constraints. The STSP-TW model enforces time windows but omits pickup-and-delivery relationships. This can be seen in the irregularity of the node visitation order, although all nodes are served within their respective time windows, there is no guarantee that pickups precede their associated deliveries. This leads to a route that may be temporally valid but infeasible from a logistics perspective.
@@ 6.4.2 Performance Analysis with a quantum solver
The objective is to assess the feasibility of expressing and solving the proposed formulations within a constrained quantum optimization framework. The experiments are intended as a proof of concept, providing insight into how the STSP-TWPD, and also STSP-PD and STSP-TW variants, can be represented and tackled through emerging hybrid quantum–classical architectures, specifically D-Wave’s LeapCQMHybrid platform. Due to the limited computational capabilities of current quantum hardware, we restrict the quantum experiments to instances from class A2, which are commonly used as benchmark instances in the literature. This part of the work is structured into three main phases.
First, we analyze the STSP-TWPD problem, under both the ABF and NBF formulations. Second, we consider the STSP-PD variant and, finally, we analyze the STSP-TW variant. All tables share the same structure. Each row corresponds to a specific problem size, indicated by $V$. The columns are organized into two main groups, comparing the performance with and without the use of AFGR. For both configurations, the reported metrics include: the average and standard deviation of the objective function value (OF Avg and OF Std), the percentage of runs leading to a feasible or optimal solution over ten repetitions per instance (% Solved), as well as the average and standard deviation of computation time (Time Avg and Time Std).
We have conducted 10 tests for each instance in order to evaluate the performance and robustness of the model. In all experiments presented in this section, the solver time limit is fixed to 10 seconds for every instance and configuration. Furthermore, whenever no feasible solution is obtained in any of the 10 runs, the corresponding entries in the table are marked with a dash “–” to indicate the absence of valid results.
@@
The experimental results are reported in E.1 (Table 46) indicate that the AFGR technique contributes to improving the scalability and numerical stability of both the ABF and NBF formulations for the STSP-TWPD.
@@
For the ABF formulation, AFGR consistently improves the solution quality and feasibility across all instance sizes. On small instances ($V=4$ and $V=5$), both configurations solve 100% of the problems; however, AFGR provides a slightly better average GAP and more stable solution times. As the problem size increases ($V\geq 6$), the benefit of AFGR becomes even more pronounced: without AFGR, no feasible solutions are found for $V=7$ to $V=10$, while with AFGR, feasible solutions are obtained for $V=7$, $V=8$, and $V=9$, with solving rates up to 20%. Notably, the GAP with AFGR remains close to or slightly lower than in the baseline formulation when a solution is available (e.g., 46–48% for V=6V=6V=9V=9), whereas the baseline often fails entirely. Additionally, average solution times are comparable or even lower with AFGR for larger instances (e.g., $V=9$: 15,32s vs 20,49s), despite the increased complexity of the feasible repair mechanism. This demonstrates that AFGR enhances both efficiency and effectiveness.
@@
Similarly, the NBF formulation benefits from AFGR across the full range of tested sizes. The results for the STSP-TWPD solved using the NBF formulation, both with and without the AFGR method. While both variants perform similarly for small instances ($V=4$ to $V=6$), AFGR slightly reduces the optimality gap (e.g., 47% vs 51% for $V=6$), improves the percentage of solved instances (e.g., 100% vs 90% for $V=5$), and maintains acceptable solution times. For larger sizes ($V\geq 7$), the baseline fails to find feasible solutions, whereas AFGR achieves partial solving success (e.g., 20–30% for V=7V=7V=9V=9) with significantly improved GAP values (e.g., 38–48%).
@@
This part of the work analyzes STSP-PD problem. At this stage, we focus solely on the ABF formulation, rather than testing both ABF and NBF, as their equivalence has already been established in subsection (4.3.1).
The results presented in E.2 (Table 47) suggest that the AFGR technique provides computational advantages, particularly in terms of model reduction and improved solver tractability. The introduction of the AFGR module leads to a systematic improvement in the average objective function value, with an average reduction in the optimality gap ranging between 2% and 10% across the tested instances. For instance, in the case with $V=6$, the gap decreases from 5% without AFGR to 2% with AFGR, highlighting the ability of AFGR to guide the search process toward higher-quality solutions. In terms of solution stability, the standard deviations of the objective values decrease significantly when AFGR is activated. This indicates greater robustness of the algorithm with respect to the inherent randomness of the heuristic optimization process. Although the addition of the AFGR mechanism introduces a slight increase in the average computational time, the total runtime remains substantially lower than that of exact approaches and is largely justified by the quality gains obtained. A more detailed analysis confirms that AFGR consistently reduces the average optimality gap across all problem sizes. Specifically, for $V=4$, the gap drops from 4% without AFGR to 1% with AFGR. For $V=6$, it decreases from 5% to 2%, and for $V=10$, a particularly challenging case, the improvement is even more pronounced, from 14% without AFGR to 6% with AFGR. Overall, the AFGR mechanism yields a gap reduction of approximately 40% to 70% compared to the baseline method, resulting in solutions that are significantly closer to the known optimum. Moreover, the percentage of successfully solved instances improves with the inclusion of AFGR, which confirms the greater efficiency of the adaptive repair mechanism in handling infeasible or borderline-feasible configurations. Integrating the AFGR module into the model without time windows leads to a substantial improvement in solution quality, enhanced robustness, and overall increased reliability of the heuristic procedure, while keeping the computational effort within acceptable limits.